# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ E.2.1 The prediction horizon $W$
In this setting, we aim to demonstrate the effect of the prediction horizon length, $W$, on the tracking performance of PLOT. We consider a static target with a $V_{T}=0$ path length. A simple target respecting this condition is the circular target, as introduced in Example B.2 with $\theta=0.06$ radians. With this target revealed and measured online, as described in Section 2, we run the PLOT algorithm repeatedly for varying number of horizon lengths. For all, we set the same initial state of $p=[0.6,0.0,0.4]^{\top}$ and forgetting factor $\gamma=0.8$. We define the augmented matrix $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{1}_{6}$ and $P_{j|j-k}=10^{-5}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$. Here, $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of all ones.
@@
E.2.2 The forgetting factor $\gamma$
@@
Figure 8: Trajectory plot of a spiral with a $V_{T}=\mathcal{O}(\sqrt{T})$ tracked with PLOT with $W=5$ and a range of values for $\gamma$, simulated for $T=2,3,5$ and $T=7$ seconds.
To show this on an example, we first fix a dynamic reference trajectory with a path length $V_{T}=\mathcal{O}(\sqrt{T})$ and perform $6$ different runs of PLOT for a range of values for $\gamma=\gamma_{a}$, with $a=0.10,0.20,0.25,0.30,0.50$ and $1.0$, and a fixed $W=5$. Note that for this path length the optimal value for $a$, suggested by Corollary 4.2 is $a=0.25$. The dynamics for the reference are generated by starting with the circle static dynamics in the previous example and increasing the radius of the circle every $\sqrt{T}$ time steps by a factor sampled uniformly from $0.7$ and $1.5$. In addition, a shift of $[-0.1,0.1]^{\top}$ is applied to $[p_{x},p_{y}]^{\top}$ at the same time step. As in the previous example, we define the augmented matrix $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2. To highlight the effect of the forgetting factor, for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{0}_{6\times 6},v_{j+k|j}=\mathbf{1}_{6}$ and $P_{j|j-k}=\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
@@
Figure 10: Regret of PLOT for the spiral target with $V_{T}=\sqrt{T}$ for a range of $\gamma$-s. $\gamma=0.8$ is chosen as per Corollary 4.2.
The dynamic target trajectory together with the system trajectory under the PLOT controller with different forgetting factors is plotted in Figure 8 for $T=2,4,9$ and $T=15$ seconds. A longer trajectory for $T=40$ seconds is provided in Figure 10 with the corresponding regret plot in Figure 10 for $T=200$ seconds and with more values of $\gamma$. It can be observed from the figures that $\gamma_{0.25}$ tuned optimally to scale with $1-T^{-0.25}$ achieves the lowest regret at the end of the horizon compared to other $a$-s. Intuitively, forgetting helps to adapt quickly to the fast-changing dynamics of the reference and lose the influence of the random initialization of the predictions exponentially fast. This can be observed on the plot, by noticing that the regret for all $\gamma$-s for the first $100$ time steps is lower than that of $\gamma=1.0$, which corresponds to PLOT without forgetting. At the other extreme, forgetting too much, as in this case with $\gamma=0.37$ hinders the learner from estimating the dynamics, just enough to cause higher regret compared to that of no forgetting with $\gamma=1$. The random initialization of learners, as well as the fast-changing dynamics of the reference, make the no forgetting tracker maintain a larger tracking error. To show that the scaling constant is independent of horizon $T$, we fix, $c_{\gamma}$ to a tuned value of $1.34$ and perform experiments with varying horizon lengths of $T=50,100,150$ and $T$. For each $T$, PLOT is run $4$ times, each time with a different $\gamma_{a}$. The regret at the end of each horizon is calculated, and the results are visualized in Figure 11. As expected from Corollary 4.2, PLOT with forgetting factor $\gamma_{0.25}$ performs the best for such a reference, across different horizon lengths.
@@
E.3 Comparison with Benchmarks
@@
We compare PLOT to 5 algorithms, namely, the Follow the Leader (FTL) algorithm Abbasi-Yadkori et al. [2014], Disturbance Action Policy controller Agarwal et al. [2019a], Riccatitron Foster and Simchowitz [2020], SS-OGD Karapetyan et al. [2023], and Naive LQR that only performs a static feedback on the currently observed tracking error. Apart from the latter, we tune each controller’s hyperparameters to achieve the best performance for the given reference, by performing a grid search over the hyperparameters. All controllers start from the same initial state and follow the same target revealed at the same point in time. We provide the details for each below in Appendix E.3.1.
@@
(d) DAP Agarwal et al. [2019a]
(e) Riccatitron Foster and Simchowitz [2020]
(f) SS-OGD Karapetyan et al. [2023]
@@
The trajectories of the target and the system under the considered algorithms are plotted in Figure 12 for the first $30$ seconds of the simulation. The corresponding dynamic regret plot for the full $T=200$ seconds horizon is shown in Figure 13. Given the dynamic nature of the reference with a sublinear path length, PLOT outperforms all of the other considered benchmarks. There are two main possible reasons for this. Firstly, our approach is aimed at dynamic targets, i.e. unlike the others, it incorporates a forgetting factor which equips it with adaptive capabilities, and enjoys dynamic regret guarantees. This allows it to deal better with such time varying references to which it is hard or impossible to fit a time invariant model better compared to the other static methods. Secondly, for such targets, the indirect approach of PLOT, i.e., learning the target dynamics and then incorporating these in the control action, is more advantageous than a direct approach deployed by all the other benchmarks. While the affine term of the optimal control action changes drastically throughout the horizon due to the sign and direction change of the target, the dynamics of the exosystem do not change as drastically. Thus, PLOT, learning the dynamics of the exosystem with forgetting is able to provide a better tracking performance, while DAP, Riccatitron, and SS-OGD learn a much smaller affine term for the controller. This results in their performance being indistinguishable from that of the naive LQR. The FTL algorithm aiming for an average best performance instead learns a control action that keeps it roughly in the center of the evolving spiral target ending up with a dynamic regret of an order higher compared to the others.
@@
Table 1 compares the average computational time for each iteration of the algorithm in milliseconds, the size of the memory that needs to be reserved and updated for the online execution of the algorithm, and the accumulated regret at the end of the given target trajectory. The average computational time is obtained by averaging the the time it takes to compute the control action at each timestep over the entire horizon of $2000$ timesteps. The memory for each controller shows the size of the past state and control variables that need to be stored in the memory. For example for the naive LQR controller, this is 00, while for SS-OGD it is $4$, as at each timestep it requires the past control input.
Naive LQR, performing only a single matrix multiplication at each timestep is unsurprisingly the fastest, requiring no extra memory. The computational time of FTL and SS-OGD is of the same order, however for this benchmark both incur a higher cost and therefore a higher regret. PLOT, Riccatitron and DAP have similar computation time, but are all an order higher compared to the Naive LQR. While compared to it, Riccatitron and DAP attain almost the same Regret, PLOT outperforms both while also requiring less memory than Riccatitron and being around a millisecond faster than DAP.
@@
PLOT (Ours) $2.10$ $45$ $\mathbf{1795}$
Riccatitron Foster and Simchowitz [2020] $2.30$ $90$ 24142414{2414}2414
DAP Agarwal et al. [2019a] $3.50$ $45$ $2454$
FTL Abbasi-Yadkori et al. [2014] 0.400.40{0.40}0.40 $\mathbf{0}$ $40004$
SS-OGD Karapetyan et al. [2023] $0.33$ $4$ $2608$
@@
E.3.1 Benchmark Implementation Details
PLOT: The prediction horizon is set to $W=5$, forgetting factor to $\gamma=0.9$ and the projection of the learnt matrices is disabled. The the augmented matrix is defined as $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{1}_{6}$ and $P_{j|j-k}=10^{-5}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
Riccatitron: Following the notation of Simchowitz [2020], the horizon of the learners is set to $h=5$ and the disturbance action policy length to $m=5$ decided upon by a search over the best parameter. The latter choice increases the memory requirement of Riccatitron in Table 1 to achieve a good tracking performance. We disable the projection, but note that even with projection enabled the results are almost identical. The online newton step (ONS) update is chosen, with The learning rate set to $\eta_{ons}=0.2$ and $\varepsilon_{ons}=1$.
DAP: As per the notation of Agarwal et al. [2019a], and to match the memory allocation for PLOT, we fix the horizon length for the disturbance action policy to $H=5$. The projection is performed as detailed in Agarwal et al. [2019a], and the learning rate is set to $\frac{0.1}{\sqrt{T}}$.
FTL: The algorithm is implemented as detailed in Abbasi-Yadkori et al. [2014] and contains no hyperparameters to be tuned.
SS-OGD: The affine control term is initialized with a vector of zeros, and the learning rate is set to $\alpha=0.001*\mathbf{I}_{4}$, following the notation of Karapetyan et al. [2023].
Naive LQR: The naive LQR controller performs a state feedback on the observed error
@@ E.3.1 Benchmark Implementation Details
Appendix F Implementation on Quadrotors

### v3

@@ F.2.1 The prediction horizon $W$
In this setting, we aim to demonstrate the effect of the prediction horizon length, $W$, on the tracking performance of PLOT. We consider a static target with a $V_{T}=0$ path length. A simple target respecting this condition is the circular target, as introduced in Example C.2 with $\theta=0.06$ radians. With this target revealed and measured online, as described in Section 2, we run the PLOT algorithm repeatedly for varying numbers of horizon lengths. For all, we set the same initial state of $p=[0.6,0.0,0.4]^{\top}$, $M=10$, and forgetting factor $\gamma=0.8$. We define the augmented matrix $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{0}_{6}$ and $P_{j|j-k}=10^{-4}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$. Here, $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of all ones.
@@
F.2.2 The forgetting factor $\gamma$
@@
Figure 11: Trajectory plot of a spiral with a $V_{T}=\mathcal{O}(\sqrt{T})$ tracked with PLOT with $W=5$ and a range of values for $\gamma$, simulated for $T=2,4,9$ and $T=15$ seconds.
To show this on an example, we first fix a dynamic reference trajectory with a path length $V_{T}=\mathcal{O}(\sqrt{T})$ and perform $5$ different runs of PLOT for a range of values for $\gamma=\gamma_{a}$, with $a=0.10,0.20,0.25,0.40,0.50$ and $1.0$. We fix $M=1$ and $W=5$. Note that for this path length the optimal value for $a$, suggested by Corollary 4.2 is $a=0.25$. The dynamics for the reference are generated by starting with the circle static dynamics in the previous example and increasing the radius of the circle every $\sqrt{T}$ time steps by a factor sampled uniformly from $0.7$ and $1.5$. In addition, a shift of $[-0.1,0.1]^{\top}$ is applied to $[p_{x},p_{y}]^{\top}$ at the same time step. As in the previous example, we define the augmented matrix $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2. To highlight the effect of the forgetting factor, for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{0}_{6\times 6},v_{j+k|j}=\mathbf{0}_{6}$ and $P_{j|j-k}=\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
@@
Figure 13: Regret of PLOT for the spiral target with $V_{T}=\sqrt{T}$ for a range of $\gamma$-s. $\gamma=0.78$ is chosen as per Corollary 4.2.
The dynamic target trajectory together with the system trajectory under the PLOT controller with different forgetting factors is plotted in Figure 11 for $T=2,4,9$ and $T=15$ seconds. A longer trajectory for $T=40$ seconds is provided in Figure 13 with the corresponding regret plot in Figure 13 for $T=200$ seconds and with more values of $\gamma$. It can be observed from the figures that $\gamma_{0.25}$ tuned optimally to scale with $1-T^{-0.25}$ achieves the lowest regret at the end of the horizon compared to other $a$-s. Intuitively, forgetting helps to adapt quickly to the fast-changing dynamics of the reference and lose the influence of the random initialization of the predictions exponentially fast. This can be observed on the plot, by noticing that the regret for $\gamma=1.0$, corresponding to PLOT without any forgetting is the highest, as expected from the theoretical results. The random initialization of learners, as well as the fast-changing dynamics of the reference, make the no-forgetting tracker maintain a larger tracking error. At the other extreme, forgetting too much, as in this case with $\gamma=0.3$ hinders the learner from estimating the dynamics, enough to cause higher regret.
To show that the scaling constant is independent of horizon $T$, we fix, $c_{\gamma}$ to a tuned value of $1.5$ and perform experiments with varying horizon lengths from $T=150$ to $T=300$ seconds. For each $T$, PLOT is run $4$ times, each time with a different $\gamma_{a}$. The regret at the end of each horizon is calculated, and the results are visualized in Figure 14. As expected from Corollary 4.2, PLOT with forgetting factor $\gamma_{0.25}$ performs the best for such a reference, across different horizon lengths.
@@
F.3 Comparison with Benchmarks
@@
We compare PLOT to 5 algorithms, namely, the Follow the Leader (FTL) algorithm (Abbasi-Yadkori et al., 2014), Disturbance Action Policy controller (Agarwal et al., 2019a), Riccatitron (Foster & Simchowitz, 2020), SS-OGD (Karapetyan et al., 2023), and Naive LQR that only performs a static feedback on the currently observed tracking error. Apart from the latter, we tune each controller’s hyperparameters to achieve the best performance for the given reference, by performing a grid search over the hyperparameters. All controllers start from the same initial state and follow the same target revealed at the same point in time. We provide the details for each below in Appendix F.3.1.
@@
(d) DAP (Agarwal et al., 2019a)
(e) Riccatitron (Foster & Simchowitz, 2020)
(f) SS-OGD (Karapetyan et al., 2023)
@@
The trajectories of the target and the system under the considered algorithms are plotted in Figure 15 for the first $30$ seconds of the simulation. The corresponding dynamic regret plot for the full $T=200$ seconds horizon is shown in Figure 16. Given the dynamic nature of the reference with a sublinear path length, PLOT outperforms all of the other considered benchmarks. There are two main possible reasons for this. Firstly, our approach is aimed at dynamic targets, i.e. unlike the others, it incorporates a forgetting factor that equips it with adaptive capabilities and enjoys dynamic regret guarantees. This allows it to deal better with such time-varying references to which it is hard or impossible to fit a time-invariant model better compared to the other static methods. Secondly, for such targets, the indirect approach of PLOT, i.e., learning the target dynamics and then incorporating these in the control action, is more advantageous than a direct approach deployed by all the other benchmarks. While the affine term of the optimal control action changes drastically throughout the horizon due to the sign and direction change of the target, the dynamics of the exosystem do not change as drastically. Thus, PLOT, learning the dynamics of the exosystem with forgetting is able to provide a better tracking performance, while DAP, Riccatitron, and SS-OGD learn a much smaller affine term for the controller. This results in their performance being indistinguishable from that of the naive LQR. The FTL algorithm aiming for an average best performance instead learns a control action that keeps it roughly in the center of the evolving spiral target ending up with a dynamic regret of an order higher compared to the others.
@@
Table 1 compares the average computational time for each iteration of the algorithm in milliseconds, the size of the memory that needs to be reserved and updated for the online execution of the algorithm, and the accumulated regret at the end of the given target trajectory. The average computational time is obtained by averaging the time it takes to compute the control action at each timestep over the entire horizon of $2000$ timesteps. The memory for each controller shows the size of the past state and control variables that need to be stored in the memory. For example for the naive LQR controller, this is 00, while for SS-OGD it is $4$, as at each timestep it requires the past control input.
Naive LQR, performing only a single matrix multiplication at each timestep is unsurprisingly the fastest, requiring no extra memory. The computational time of FTL and SS-OGD is of the same order, however, for this benchmark, both incur a higher cost and therefore a higher regret. When PLOT is implemented by approximating step $8$ of Algorithm 1 with an unweighted Frobenious norm projection, it can be implemented by a thresholding operation on singular values, requiring no quadratic programs to be solved. In this case, PLOT, Riccatitron, and DAP have similar computation times but are all an order higher compared to the Naive LQR. While compared to it, Riccatitron and DAP attain almost the same regret, PLOT outperforms both, requires less memory than Riccatitron, and is around a millisecond faster than DAP. When the full weighted projection is carried out in Algorithm 1, step $8$, by approximately solving it with a quadratic program consuming $9.3$ milliseconds, PLOT attains the lowest regret.
@@
PLOT (Ours) $10.7^{*}$ $54$ $\mathbf{1661}$
PLOT (Ours, w/ unweighted projection) $2.10$ $54$ 16851685{1685}1685
Riccatitron (Foster & Simchowitz, 2020) $2.30$ $90$ 24142414{2414}2414
DAP (Agarwal et al., 2019a) $3.50$ $45$ $2454$
FTL (Abbasi-Yadkori et al., 2014) 0.400.40{0.40}0.40 $\mathbf{0}$ $40004$
SS-OGD (Karapetyan et al., 2023) $0.33$ $4$ $2608$
@@
* - 9.3 ms spent on the weighted projection with a QP solver.
F.3.1 Benchmark Implementation Details
PLOT: The prediction horizon is set to $W=6$, the forgetting factor to $\gamma=0.9$, and $M=10$. The the augmented matrix is defined as $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{0}_{6}$ and $P_{j|j-k}=10^{-4}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
Riccatitron: Following the notation of (Simchowitz, 2020), the horizon of the learners is set to $h=5$ and the disturbance action policy length to $m=5$ decided upon by a search over the best parameter. The latter choice increases the memory requirement of Riccatitron in Table 1 to achieve a good tracking performance. We disable the projection, but note that even with projection enabled the results are almost identical. The Online Newton Step (ONS) update is chosen, with The learning rate set to $\eta_{ons}=0.2$ and $\varepsilon_{ons}=1$.
DAP: As per the notation of (Agarwal et al., 2019a), and to match the memory allocation for PLOT, we fix the horizon length for the disturbance action policy to $H=5$. The projection is performed as detailed in (Agarwal et al., 2019a), and the learning rate is set to $\frac{0.1}{\sqrt{T}}$.
FTL: The algorithm is implemented as detailed in (Abbasi-Yadkori et al., 2014) and contains no hyperparameters to be tuned.
SS-OGD: The affine control term is initialized with a vector of zeros, and the learning rate is set to $\alpha=0.001*\mathbf{I}_{4}$, following the notation of (Karapetyan et al., 2023).
Naive LQR: The naive LQR controller performs state feedback on the observed error
@@ F.3.1 Benchmark Implementation Details
F.3.2 Random Targets
The dependence of the dynamic regret of PLOT on $V_{T}$ implies that for targets having a dynamic structure, including the one in the preceding example, PLOT’s tracking performance is better compared to targets that lack dynamics. While the latter case is still covered by the structure imposed on the targets, as discussed in Section C, the regret can scale linearly with $T$ in the worst case when $r_{t}$ at each time step is independent from the others. In this case, the indirect approach of PLOT may result in a poorer performance compared to direct approaches like DAP, Riccatitron, or SS-OGD when $\gamma$, $W$, and $M$ are not optimally tuned for the specific realization.
Figure 17: Average Dynamic Regret of Online Control algorithms applied to the online tracking problem of the unknown target following (31).
To showcase this marginal example, we consider random targets sampled from the uniform distribution
$r_{t}\sim\mathcal{U}\left(\begin{bmatrix}-1\\ -1\end{bmatrix},\begin{bmatrix}1\\ 1\end{bmatrix}\right).$ (31)
We compare PLOT with the same benchmark controllers whose tunable parameters are the same as in Section F.3.1. For PLOT we take $W=2$, $\gamma=0.02$, and $M=0.7$, leaving the other parameters unchanged from Section F.3.1.
We perform $10$ independent realizations of the random targets over a horizon of length $T=150$ seconds and compute the averaged regret over these runs for each of the six considered algorithms, shown in Figure 17. The random reference updates combined with the low forgetting factor $\gamma$ result in $\hat{S}_{t}$-s with large singular values, and, consequently, large inputs, while the optimal action in this case is staying close to the origin. While the performance can be improved significantly by lowering $M$ and increasing $\gamma$, bringing the averaged regret to around $25$, this example shows that for a completely random agent with no dynamics a direct approach may perform better than an indirect one like PLOT.
F.4 A Naive-RLS Controller
The multi-predictor setup of PLOT is crucial for its stability and regret performance. To showcase this we consider also a simpler controller that we refer to as Naive-RLS in Algorithm 4. In contrast to PLOT, it obtains $W$-step ahead dynamics predictions by recursively applying only the one step ahead estimate $\hat{S}_{t+1|t}$, resulting in $r_{t+1}=\hat{S}_{t+1|t}r_{t},~{}r_{t+2|t}=\hat{S}^{2}_{t+1|t}r_{t},~{}\ldots,~{}r_{t+W|t}=\hat{S}^{W}_{t+1|t}r_{t}$, for $p=1$. As detailed in Section 3, this can result in unstable predictions in particular when $S_{t}$ is time-varying. For $p>1$, we define $z_{t|t}=z_{t}$, $r_{t|t}=r_{t}$, and, for $k=1,\ldots,W-1$
$z_{t+k|t}=\begin{cases}\begin{bmatrix}r^{\top}_{t+k|t}&\cdots&r_{t|t}&r_{t-1}\cdots&r^{\top}_{t-p+k+1}\end{bmatrix}^{\top}\;\text{if}\;p>k,\\ \begin{bmatrix}r^{\top}_{t+k|t}&\cdots&r^{\top}_{t-p+k+1|t}\end{bmatrix}^{\top}\;\text{if}\;p\leq k.\par\end{cases}$
Algorithm 4 Naive-RLS
0:  Horizon $W$, forgetting factor $\gamma\in(0,1)$
1:  Compute $X,K,\{K_{0},...,K_{W-1}\}$ as in (6), (8), (9).
2:  Initialize $1$ RLS learner according to the Algorithm 1, for $k=1$.
3:  for t = 0, …, $T-1$ do
4:     Observe system state $x_{t}$, target state $r_{t}$.
5:     for $k=1,...,W$ do
6:        if $t+k\leq T$ then
7:           Update the $1$-step-ahead predictor and obtain $\hat{S}_{t+1|t}$ according to the Algorithm 1.
8:           Predict $r_{t+k|t}=\hat{S}_{t+1|t}z_{t+k-1|t}$
9:        else
10:           Set $r_{t+k-1|t}=0$.
11:        end if
12:     end for
13:     Compute $u_{t}^{\pi}$ as in (13)
14:  end for
To demonstrate the failure of Naive-RLS for agents with time-varying dynamics, we consider the static (circular) and dynamic references of Sections F.2.1 and F.2.2, respectively. Figure 19 depicts the former case, showing the comparable performance of Naive-RLS to PLOT, since the static $S$ remains stable even when exponentiated. In the latter case with a dynamic $S_{t}$, shown in Figure 19, Naive-RLS becomes unstable due to exploding matrices $\hat{S}_{t+k|t}$. The implementation details of both controllers are provided below.
PLOT: The prediction horizon is set to $W=10$, the forgetting factor to $\gamma=0.8$, and $M=1$. The the augmented matrix is defined as $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{0}_{6}$ and $P_{j|j-k}=10^{-4}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
Naive-RLS: The prediction horizon is set to $W=10$, the forgetting factor to $\gamma=0.8$, and $M=1$. The the augmented matrix is defined as $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ for $t=1,\ldots,T-1$ as in Section 2, and for all $k=1,\ldots,W$, the $k$ learners are initialized as follows: $\hat{S}_{j+k|j}=\mathbf{I}_{6},v_{j+k|j}=\mathbf{0}_{6}$ and $P_{j|j-k}=10^{-4}\times\mathbf{I}_{7}$, for all $j=-1,\ldots,k-2$.
Figure 18: Naive-RLS shows comparable performance to PLOT on the static, circular target from Example C.2.
Figure 19: Naive-RLS becomes unstable while PLOT manages tracks the dynamic target of Section F.2.2.
Appendix G Implementation on Quadrotors